OpenStudy (anonymous):

I need help putting x+2y=4 into y intercept form

OpenStudy (anonymous):

2y=4-x \[y=\frac{4-x}{2}\]

OpenStudy (anonymous):

y=-1/2x+2

OpenStudy (anonymous):

\[y=-\frac{1}{2}(x-4)\]

OpenStudy (mathstudent55):

Slope intercept form is y = mx + b You need to solve the equation for y, which is what clavoie did.

OpenStudy (anonymous):

but i dont get where he got 1/2 from though?

OpenStudy (mathstudent55):